OpenStudy (staceyg):

A cylinder with radius r and height 2r + 4 contains a cube with edge length r√2, as shown. What fraction of the cylinder's volume is taken up by the cube? Write your answer in simplified form.

OpenStudy (staceyg):

@SithsAndGiggles

OpenStudy (science0229):

Is there some kind of a diagram for this?

OpenStudy (science0229):

@staceyg

OpenStudy (staceyg):

There's a model of the cylinder and cube.

OpenStudy (staceyg):

@science0229

OpenStudy (science0229):

Sorry. Seems like we don't need a diagram for this one. Do you know how to find the volume of a cylinder and a cube?

OpenStudy (science0229):

Volume of a cylinder=\[\pi r^2h\] Volume of a cube=\[s^3\]

OpenStudy (science0229):

Cylinder=\[\pi r^2(2r+4)\] Cube=\[(r \sqrt{2})^3=2\sqrt{2}r^3\]

OpenStudy (science0229):

fraction of the cylinder's volume is taken up by the cube=\[\frac{ cube }{ cylinder }=\frac{ \pi r^2(2r+4) }{ 2\sqrt{2}r^3 }=\frac{ \pi(r+2) }{ \sqrt{2}r }=\frac{ \pi \sqrt{2}(r+2) }{ 2r }\]

OpenStudy (science0229):

This is the simplest that you can get

OpenStudy (science0229):

Any questions?

OpenStudy (science0229):

@staceyg

OpenStudy (staceyg):

No, thanks so much for explaining :)